OpenStudy (cooltowl):

Solve for x: (x−4)^2−9=0

OpenStudy (anonymous):

what is (x-4)(x-4)?

OpenStudy (tkhunny):

What you should ALWAYS do is FACTOR. A^2 - 9 = 0 is solved this way. (A+3)(A-3)=0 B^2 - 9 = 0 is solved this way. (B+3)(B-3)=0 Frog^2 - 9 = 0 is solved this way. (Frog+3)(Frog-3)=0 How then, do you suppose, we should solve: (x-4)^2 - 9 = 0?

OpenStudy (anonymous):

(x-4)^2-3^2......solve for (a+b)(a-b)

OpenStudy (anonymous):

lol